OpenStudy (anonymous):

Every evening Jenna empties her pockets and her change in a jar. At the end of the week she counts all her money. One week she had 50 coins, all of them dimes and quarters. When she added them up, she had a total of $10.25. How many dimes did she have?

OpenStudy (luckycoins888):

let d be the number of dimes let q be the number of quarters d + q = 40------> multiply this by -25 10d + 25q = 775 -25d -25q = -1000 10d +25q = 775 ----------------------------- -15d = -225 d = 15 dimes @atlantaslast There you go XD